OpenStudy (anonymous):

please help i am truly about to cry, any help just please help please The table below shows the number of hours high school students of different cities of two states spend on indoor sports each week: State A 7 9 5 4 25 21 6 6 8 State B 14 12 10 13 15 14 1 15 16 Part A: Why is the histogram for State A different in shape, center, and spread than the histogram or State B? (5 points) Part B: In which state did the students spend more average time on indoor sports? Justify your answer. (5 points)

OpenStudy (anonymous):

please help i will definitely reward you with a medal

OpenStudy (mokeira):

for part B find the mean for both and compare

OpenStudy (anonymous):

how do i find the mean

OpenStudy (anonymous):

add up all the numbers and then divide it by how many numbers there are

OpenStudy (anonymous):

so add them up and then divide both by 9

OpenStudy (mokeira):

State A \[\frac{ (7+9+5+4+25+21+6+6+8) }{9 }=\frac{ 91 }{ 9 }=simplify\]

OpenStudy (anonymous):

state A= 10.1?

OpenStudy (mokeira):

State B \[\frac{ (14+12+10+13+15+14+1+15+16) }{ 9 }=\frac{ 110 }{ 9 }\]= simplify

OpenStudy (anonymous):

and state b: 12.2?

OpenStudy (dumbcow):

so dramatic ...

OpenStudy (mokeira):

yes @zakeya

OpenStudy (mokeira):

so which is more

OpenStudy (anonymous):

yes drama queen over here @dumbcow

OpenStudy (anonymous):

state b is more?

OpenStudy (mokeira):

correct @zakeya

OpenStudy (anonymous):

how about part a

OpenStudy (anonymous):

@Mokeira

OpenStudy (mokeira):

different in shape because they have different values

OpenStudy (mokeira):

Different in everything because they have different values

OpenStudy (mokeira):

When the values are different, the mean will be different, the mode will be different, the width will be different, everything will be diefferent

OpenStudy (mokeira):

@zakeya do you get the point of what I mean?

OpenStudy (anonymous):

yes thankyou so much

OpenStudy (mokeira):

you welcome :)

OpenStudy (anonymous):

do you mind helping me with one more?

OpenStudy (mokeira):

no problem @zakeya